OpenStudy (anonymous):

solve = e^x ln(x + 1/2) = 0 i know the solution is 1/2 but i have no idea how they got there. i keep getting 1/4. what i did was this: e^x ln(x + 1/2)= log e (x) . Log e ( x+ 1/2)= Log e (2x + 1/2) (2x + 1/2) = 1 (because e^1= 0) X= 1/4

OpenStudy (aum):

e^x * ln(x + 1/2) = 0 Either e^x = 0 OR ln(x+1/2) = 0 OR Both = 0 e^x approaches zero as x-> - infinity but never attains zero. So ln(x+1/2) = 0 which implies x + 1/2 = 1 or x = 1/2 (because log(1) = 0)

OpenStudy (anonymous):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)